OpenStudy (anonymous):

Solve for x: the quantity of x plus 4 over 3 = 2

OpenStudy (anonymous):

Yes I can thank you

OpenStudy (anonymous):

\[ \frac {x+4}3=2\\ x+4 =6 \]

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

I got the answer but thank you any way hun